OpenStudy (anonymous):

Given f(x)=x^3-6x^2+9x and g(x)=4. Find the coordinates of the points common to the graphs of f and g. -find all zeros of f -if the domain of f is limited to the closed interval [0,2], what is the range of f?

OpenStudy (anonymous):

#1 solve y=x^3-6x^2+9x and y=4 that is 4=x^3-6x^2+9x These are the x coordinates... The y coordinate will be 4 (as g(x)=4

OpenStudy (anonymous):

#2 solve for f(x)=0 to find the zeros of f

OpenStudy (anonymous):

So,the y value is a given of Four? And what do i do for tee zeros? *the